i'm sorry, i left ouy sq 2.the question again was solve the quadratic formula for 6x^2+3x-18=0. help
Answer by algebrapro18(249)

I use the master product method because its easier than guess and check.

6x^2+3x-18=0
3(2x^2+x-6)=0

12|
------
1 | 12
2 | 6
3 | 4

(2x^2+4x-3x-6)=0
(2x(x+2)-3(x+2))=0
(2x-3)(x+2)=0

2x-3=0 x+2=0
2x=3 x=-2
x=3/2

so the two solutions are x = 3/2 and x = -2